OpenStudy (oleg3321):

Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15.

OpenStudy (oleg3321):

If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. Marco found five charms that he likes. Using your function, determine the cost of the bracelet he would make. As they walked down the hall of the mall, Marco saw this sales flyer for Store B on a bench.

OpenStudy (oleg3321):

@SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman @SolomonZelman

OpenStudy (solomonzelman):

so we can say f(x)=15x+85 where x is a number of charms I would just graph the function, and to know what could they buy if they have 250$ I would plug in 250 for f(x) and solve for x.

OpenStudy (oleg3321):

so then how many charms could they could purchase with a bracelet if they had saved between $85 and $250

OpenStudy (oleg3321):

@samjordon

OpenStudy (oleg3321):

@tkhunny

OpenStudy (anonymous):

Ok the question is asking you to graph the function with a range of \( 85 \leq f(x) \leq 250\) Since this is a linear equation as solomon has demonstrated above all you seemingly would have to do if plug in f(x)=85 and determine what x equals and do the same for f(x)=250 and then draw a straight line connecting both of them. At least thats what I would do

OpenStudy (solomonzelman):

yes, apparently so. It would be a segment.

OpenStudy (anonymous):

The second part is asking you to find how much it would cost if you bought the bracelet and 5 charms. We would use Solomons equation where x represents how many charms purchased and f(x) represents the cost of the bracelet. We would plug in x=5 and solve to find out what our total cost would be f(5)=15(5)+85 =160 Hence we conclude that purchasing a bracelet and 5 charms would cost a total of $160

OpenStudy (oleg3321):

ok.and lastly part 3?

OpenStudy (anonymous):

What is part 3???